OpenStudy (e.cociuba):

*****HELP!!*****METAL!!******What volume of each of these solutions would contain 58.44g of NaCl? A.) 0.10 M NaCl B.) 3.0 M NaCl C.) 5.5 M NaCl

OpenStudy (e.cociuba):

@robtobey, @jes230, @jim_thompson5910, @phi HELP PLZ!!

OpenStudy (e.cociuba):

@pooja195, @abc123!, @april115,@abb0t ANYONE PLZ HELP! :)

OpenStudy (e.cociuba):

@Jemurray3

OpenStudy (anonymous):

\[ \frac{\text{# of moles}}{\text{Volume}} = \text{ Concentration } = \frac{\text{Mass of Sample}}{\text{ (Molar Mass)}\cdot \text{(Volume)}}\] \[ \text{Volume} = \frac{\text{Mass of Sample}}{\text{(Molar Mass)}\cdot \text{(Concentration)}} \]

OpenStudy (e.cociuba):

Thanks So Much! This will so help me:) Bless ur heart! Thanks! :)